Consider a small unit of a factory where there are 5 employees : a supervisor and four labourers. The workers earn a salary of Rs. 5,000 per month each while the supervisor gets Rs. 15,000 per month. Calculate the mean, median and mode of the salaries.

1
Expert's answer
2022-04-12T06:54:10-0400

"Salaries: 5000,5000,5000,15000\\Mean:\\\frac{5000+5000+5000+15000}{4}=7500\\Median:\\\frac{x_{2}^{*}+x_{3}^{*}}{2}=\frac{5000+5000}{2}=5000\\Mode: the\,\,value\,\,with\,\,\max imum\,\,frequency, that\,\,is\,\,5000"